INAF queue observing, 2014 Oct 28-29

Observer: E. Sani, R. Carini
Telescope Operator: S. Allanson
Support Astronomer: D. Thompson on call
Instrument: MODS, LBC-PARALLEL

Summary

When the Dome is opened the humidity is 76%, no wind
Variable seeing at the beginning of the scientific operations.
ModsDisp works well.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Standard feige110 dual and red grating   airmass 1.5
Fusi Pecci S H7 dual grating 1h COMPLETED
Fusi Pecci T G002 dual grating 45 m COMPLETED
Saracco F XLLSM1 red grating + LBCred 2 h 45 m  
Israel A1 catsj021950 dual grating 15 m  
Standard star SA98   airmass 1.23

Details

00:05 Run MODS Calib for Saracco_F program
  • grpix_r.cal mods1r.20141029.0001-0005 SlitLess flats, RED grating
  • mods.5.xllms1.cal mods1r.20141029.0008-0012 Mask5 flats, mods1r.20141029.00013-00015 Mask5 arcs.
  • images mods1r.0006-0007 are saturated mask flats, decrease the exptime from 4s to 2s.

00:51 The TO opens the Dome

01:34 acqMODS feige110. acq
  • modsAlign -r mods1r.0016.fits
  • Error: command 'offsetxy rel' timed out after 60 seconds
  • acqMODS feige110.acq
  • connection problem between MODS and the telescope, the acquisition crashed. we call Dave. TCS from MODS died, we restored it and run the acquisition again.
  • modsAlign -r mods1r.0017-0018.fits

01:55 execMODS feige110.obs mods1r.0019-21,  mods1b.0001-0003 dual grating
  • mod1r.0022-24 red grating

02:16 slew to scientific target Fusi Pecci S
  • acqMODS H7_ls.acq mods1r.0025-0027

02:31 execMODS H7_ls.obs==mods1r.0028- 30,mods1b.0004-0006 ==
  • at the beginning of the integration the seeing is 0.9
  • during the 2nd frame integration the seeing as a spike reaching 2.0'', then suddenly drops back to 0.8. The spike should not affect the data.

03:00 slew to Fusi Pecci T
  • acqMODS G002_ls.acq== mods1r.0030-32==
  • change the guide star
  • acqMODS G002_ls.acq== mods1r.0033-36==

03.06 execMODS G002_ls.obs==mods1r.0037-39, mods1b.0007-0009==

04:44 begin the procedure to the parallel mode for Saracco F program
  • first collimation on the red channel

05:02 acqMods mods.5.xllsm1_recover.acq (recovery procedure) mods1r.0040-42=

05:05 collimation

05:16 run XLSSm1_I_180s.ob

05:22 execMODS mods.5.xllsm1_3.obs mods1r.0043-0044
  • finish the 1/2 OB started the past night
  • at the beginning of scientific operations the mean FWHM is 0.7

05:50 after 9 images, collimation needed
  • some defocus in = lbcr.20131029.054549.fits lbcr.20141029.055040.fits=

05:54 Run XLSSm1_I_180s.ob
  • the seeing measure on the first I-band image is 0.8'''
  • 3 good images (to be checked)

06:04 execMODS mods.5.xllsm1_4.obs mods1r.0045-0048

06:12 After 4 images the data quality decreases (some defocus, some astigmatism) e.g. lbcr.060528.
  • This is stressed by a very good seeing.
  • dofpia, /X2, /redonly on the scientific Run

06:17 Run XLSSm1_I_180s.ob
  • image lbr.062058 is blurry, it was integrated while MODS was offsetting
  • the measured seeing is 0.6''
  • offsetting MODS, image lbr.065308 is blurry

07:03 dofpia, /X2, /redonly on the scientific field mask5

07:12 execMODS mods.5.xllsm1_5.obs mods1r.0049-0050 *Run XLSSm1_I_180s.ob *image lbcr.071404 is blurry
  • the measured seeing is 0.66

07:40 dofpia, /X2, /redonly on the scientific field mask5

08:00 MODS error message
  • the script crashed while committing the 3rd offset. Aborting the exposure and Retryin to commit the offset, the script performs a series of 2 offsets.
  • acquire images to control the offsets mods1r.0051-54 are blurry and the offsets are not good.
  • The IMCS was uncheked, that's genetared the blurry images. We need to recover the acquisition.
  • lbcr.071406, lbcr.071408, lbcr.071409, are made during the offset of mods

08:10 acqMODS mods.5.xllsm1_recover.acq (recovery procedure) mods1r.0055-56
  • dofpia, /X2, /redonly on the scientific field mask5

08:15 expecMODS mods.5.xllsm1_5.obs mods1r.0057-58

8.55 expecMODS mods.5.xllsm1_5.obs mods1r.0059

09:30 slew to Israel A1 program
  • acqMODS catsj021950_ls.acq mods1r.0060-60

10:10 execMODS catsj021950_ls.obs mods1r.0061, mods1b.0010
  • we stop the integration after 15 min to Slew to Dotto_DDT. We see the targeted spectrum in this short frame.

10:25 We ask the TO to switch to LBC Binocular

10.40 dofpia, /X2 /first on copointing star WT10-271
  • can not converge the blue

11.10 the TO clears the left mirror optic
  • dofpia, /X2 /first
  • can not converge both

11.19 *the TO clears the left and right mirror
  • dofpia, /X2 /first
  • dofpia is conviced the red channel converge properly, but it is clearly not true: it is measuring Z4>500 but does not apply any correction

11:28 dofpia, /X2
  • Red still too bad, too high forces. The TO clears the red channel

11:36 dofpia, /X2, /redonly, /first
  • dofpia is sending corrections at the Z4 even if the values are reasonable.

11:56 slew to standar star SA98
  • dofpia, /X2, /first

*12.15 * copointing with the stars ACTO 0197
  • copointing takes a lot because red rotator has to unwrap

12.25 iraf can not comunicate with ds9
  • RB_science can not comunicate with iraf

12.35 run the standard star SA98 , airmass = 1.23

12.50 We star skyflats
  • Run V- R-Bessel the last image in R and V in position angle 180 are not good

1.49 bias with MODS

-- %USERSIG{EleonoraSani - 2014-10-29}%

Comments

-- %BUBBLESIG{RobertaCarini - 2014-10-29}%
 
Topic revision: r7 - 19 Dec 2014, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback